A cash register contains only five dollar and ten dollar bills. It contains twice as many five's as ten's and the total amount of money in the cash register is 740 dollars. How many ten's are in the cash register?
Answer by checkley77(12844)

2X*5+X*10=740
10X+10X=740
20X=740
X=740/20
X=37 TEN DOLLAR BILLS.
2*37=74 FIVE DOLLAR BILLS.
PROOF:
74*5+37*10=740
370+370=740
470=740